0.4grams of a compound on complete combustion gave 56ml of co2 at760mm and 0degree centigrade.the percentage of carbon in the compound is?
1
Expert's answer
2013-12-23T09:52:29-0500
Calculate the amount of the evolved CO2:
n(CO2) = 0.056/22.4 = 0.0025 mol
m(C) = n(CO2)*M(C) = 0.03 g
Hence we can obtain the percentage of carbon:
w(C) = m(C)/m(compound) = 0.075 = 7.5%